# Federated Continual Recommendation

### v2

@@ Abstract.
The increasing emphasis on privacy in recommendation systems has led to the adoption of Federated Learning (FL) as a privacy-preserving solution, enabling collaborative training without sharing user data. While Federated Recommendation (FedRec) effectively protects privacy, existing methods struggle with non-stationary data streams, failing to maintain consistent recommendation quality over time. On the other hand, Continual Learning Recommendation (CLRec) methods address evolving user preferences but typically assume centralized data access, making them incompatible with FL constraints. To bridge this gap, we introduce Federated Continual Recommendation (FCRec), a novel task that integrates FedRec and CLRec, requiring models to learn from streaming data while preserving privacy. As a solution, we propose F3CRec, a framework designed to balance knowledge retention and adaptation under the strict constraints of FCRec. F3CRec introduces two key components: Adaptive Replay Memory on the client side, which selectively retains past preferences based on user-specific shifts, and Item-wise Temporal Mean on the server side, which integrates new knowledge while preserving prior information. Extensive experiments demonstrate that F3CRec outperforms existing approaches in maintaining recommendation quality over time in a federated environment. [^1] https://github.com/Jaehyung-Lim/F3CRec-CIKM-25
Federated Learning, Continual Learning, Recommender Sysetms
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Retrieval models and rankingccs: Information systems Recommender systemsccs: Computing methodologies Distributed artificial intelligence
@@ 4.3.1. Pre-aggregation
1:Function: ClientUpdate($u,Q_{g}^{t,r},t$)
2:Set $Q_{u}^{t,r}\leftarrow Q_{g}^{t,r}$
3:Set $\Phi_{u}^{t,r}$ with latest private parameter
4:for each epoch e=1,2,,Ee=1,2,...,E do
5:  for batch $b\in\mathcal{D}^{t}(u)$ do
6:   Compute recommendation loss $\mathcal{L}_{\text{Rec}}^{t,r}(u)$
7:   if $t>0$ then
8:     Obtain $M_{u}^{t,r}$ using Eq. 4
9:     Compute distillation loss $\mathcal{L}_{\text{KD}}^{t,r}(u)$ using Eq. 5
10:   else
11:     Set $\mathcal{L}_{\text{KD}}^{t,r}(u)\leftarrow 0$
12:   end if
13:   $\mathcal{L}_{\text{total}}^{t,r}(u)\leftarrow\mathcal{L}_{\text{Rec}}^{t,r}(u)+\lambda_{\text{KD}}\cdot\mathcal{L}_{\text{KD}}^{t,r}(u)$
14:   $Q_{u}^{t,r}\leftarrow Q_{u}^{t,r}-\eta\nabla_{Q_{u}^{t,r}}\mathcal{L}_{\text{total}}^{t,r}(u)$
15:   $\Phi_{u}^{t,r}\leftarrow\Phi_{u}^{t,r}-\eta\nabla_{\Phi_{u}^{t,r}}\mathcal{L}_{\text{total}}^{t,r}(u)$
@@
1:Initialize $Q_{g}^{0}$
2:Initialize $\Phi_{u}$ for all $u\in\mathcal{U}^{0}$
3:for each task $t=0,1,\dots,T$ do
4:  if $t>0$ then
5:   Set $Q_{g}^{t}\leftarrow Q_{g}^{t-1}$
6:   Initialize new item embeddings: $Q_{g,i}^{t}$ for $i\in\mathcal{I}^{t+1}\setminus\mathcal{I}^{t}$
7:   Initialize new users: $\Phi_{u}$ for $u\in\mathcal{U}^{t+1}\setminus\mathcal{U}^{t}$
8:  end if
9:  for each round $r=1,2,\dots,R$ do
10:   $\mathcal{U}^{t,r}$ $\leftarrow$ Server randomly selects a subset of clients
11:   for each client $u\in\mathcal{U}^{t,r}$ in parallel do
12:     $Q_{u}^{t,r}\leftarrow\texttt{ClientUpdate}(u,Q_{g}^{t,r},t)$
13:   end for
14:   Compute $Q_{g^{\prime}}^{t,r}\leftarrow\frac{1}{|\mathcal{U}^{t,r}|}\sum_{u\in\mathcal{U}^{t,r}}Q_{u}^{t,r}$
15:   if $t=0$ then
16:     Set $Q_{g}^{t,r+1}\leftarrow Q_{g^{\prime}}^{t,r}$
17:   else
18:     Obtain $Q_{g}^{t,r+1}$ using Eq.13
19:   end if
20:  end for
21:end for
@@ 5.2. Main Performance
Misalignment of existing CLRec with FedRec. F3CRec outperforms RLP2 and SPP, which adapt centralized CLRec methods to FedRec. F3CRec surpasses RLP2 by 4% to 22%. While RLP2 leverages global knowledge for modeling preference shifts, FL constraints limit it to local knowledge, leading to performance drops. In contrast, F3CRec captures preference shifts individually and item knowledge shifts globally, yielding superior performance. SPP achieves the best baseline performance in about half of the cases, yet F3CRec still outperforms SPP by 2% to 16%. Unlike SPP’s uniform retention across users and items, F3CRec uses personalized retention at the client and item-specific weighting at the server, enhancing effectiveness. F3CRec also outperforms Reg and KD, which rely solely on user-specific data, with gains of 4%–50% and 1%–33%, respectively. Reg applies a naive retention approach without considering preferences, whereas KD assumes equal retention for all users. F3CRec explicitly accounts for individual preference shifts, leading to superior results. Overall, existing CLRec methods are misaligned with FedRec due to neglecting decentralization. F3CRec effectively addresses this by retaining knowledge on both client and server sides.
@@ 5.3.2. Analysis on Adaptive Replay Memory
ther the proposed preference shift measure effectively distinguishes static users and dynamic users. We compute the shift (Eq.2) for users appearing in both $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$ after the first training round on $\mathcal{D}^{1}$. Users in the bottom 20% of preference shift are labeled static, and those in the top 20% dynamic. To assess preference change, the model trained on $\mathcal{D}^{0}$ is first evaluated on its test set of the same block. The model is then further trained on $\mathcal{D}^{1}$ and re-evaluated on the test set of $\mathcal{D}^{0}$ to assess knowledge preservation (Table 5) and performance degradation (Table 4), which is measured as $\frac{a_{0,0}-a_{1,0}}{a_{0,0}}$, where $a_{(t,s)}$ denotes performance when trained on $\mathcal{D}^{t}$ and evaluated on $\mathcal{D}^{s}$. Experiments are conducted using FedMF with the N@20 metric on the ML-100K and Lastfm-2K datasets. Table 4 and 5 show that the proposed method effectively distinguishes static and dynamic users. Static users achieve higher N@20 and lower preference degradation rates while dynamic users show the opposite. This demonstrates the effectiveness of our preference shift measure and highlights its capability for adaptive, user-specific continual learning leveraging solely users’ own data without others.
@@ 5.6. Privacy-Preserving
Although private parameters and interaction data remain private in FL settings, information leakage may occur when transmitting public parameters to the server. Given a user $u$ with $n_{u}^{+}$ positive and $n_{u}^{-}$ negative items, the server can compare $Q_{g}^{t,r-1}$ and $Q_{u}^{t,r}$ to estimate the probability of an item being positive as $n_{u}^{+}/(n_{u}^{+}+n_{u}^{-})$, compromising privacy. To mitigate this, we add zero-mean Laplace noise to user updates before transmission: $Q_{u}^{t,r}=Q_{u}^{t,r}+\text{Laplace}(0,\lambda)$, where $\lambda$ controls the noise intensity. While not a formal local differential privacy (LDP) mechanism, this strategy is loosely motivated by LDP principles (Zhang et al., 2024b, 2023; Choi et al., 2018). We evaluate the performance of F3CRec with FedMF under varying noise intensities $\lambda\in{0,0.1,0.2,0.3,0.4,0.5}$ across four datasets, averaged over all data blocks (Table 8). Results show that even with stronger noise, performance degradation remains low, indicating our method reduces privacy leakage during aggregation without sacrificing utility.

### v3

@@ Abstract.
The increasing emphasis on privacy in recommendation systems has led to the adoption of Federated Learning (FL) as a privacy-preserving solution, enabling collaborative training without sharing user data. While Federated Recommendation (FedRec) effectively protects privacy, existing methods struggle with non-stationary data streams, failing to maintain consistent recommendation quality over time. On the other hand, Continual Learning Recommendation (CLRec) methods address evolving user preferences but typically assume centralized data access, making them incompatible with FL constraints. To bridge this gap, we introduce Federated Continual Recommendation (FCRec), a novel task that integrates FedRec and CLRec, requiring models to learn from streaming data while preserving privacy. As a solution, we propose F3CRec, a framework designed to balance knowledge retention and adaptation under the strict constraints of FCRec. F3CRec introduces two key components: Adaptive Replay Memory on the client side, which selectively retains past preferences based on user-specific shifts, and Item-wise Temporal Mean on the server side, which integrates new knowledge while preserving prior information. Extensive experiments demonstrate that F3CRec outperforms existing approaches in maintaining recommendation quality over time in a federated environment. Our code is available at https://github.com/Jaehyung-Lim/F3CRec-CIKM-25.
Federated Learning, Continual Learning, Recommender Systems, Federated Continual Recommender Systems
journalyear: 2025copyright: ccconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3761268isbn: 979-8-4007-2040-6/2025/11ccs: Information systems Retrieval models and rankingccs: Information systems Recommender systemsccs: Computing methodologies Distributed artificial intelligence
@@ 4.3.1. Pre-aggregation
Function: ClientUpdate($u,Q_{g}^{t,r},t$)
1:Set $Q_{u}^{t,r}\leftarrow Q_{g}^{t,r}$
2:Set $\Phi_{u}^{t,r}$ with latest private parameter
3:for each epoch e=1,2,,Ee=1,2,...,E do
4:  for batch $b\in\mathcal{D}^{t}(u)$ do
5:   Compute recommendation loss $\mathcal{L}_{\text{Rec}}^{t,r}(u)$
6:   if $t>0$ then
7:     Obtain $M_{u}^{t,r}$ using Eq. 4
8:     Compute distillation loss $\mathcal{L}_{\text{KD}}^{t,r}(u)$ using Eq. 5
9:   else
10:     Set $\mathcal{L}_{\text{KD}}^{t,r}(u)\leftarrow 0$
11:   end if
12:   $\mathcal{L}_{\text{total}}^{t,r}(u)\leftarrow\mathcal{L}_{\text{Rec}}^{t,r}(u)+\lambda_{\text{KD}}\cdot\mathcal{L}_{\text{KD}}^{t,r}(u)$
13:   $Q_{u}^{t,r}\leftarrow Q_{u}^{t,r}-\eta\nabla_{Q_{u}^{t,r}}\mathcal{L}_{\text{total}}^{t,r}(u)$
14:   $\Phi_{u}^{t,r}\leftarrow\Phi_{u}^{t,r}-\eta\nabla_{\Phi_{u}^{t,r}}\mathcal{L}_{\text{total}}^{t,r}(u)$
15:  end for
@@
Initialize $Q_{g}^{0}$
Initialize $\Phi_{u}$ for all $u\in\mathcal{U}^{0}$
1:for each task $t=0,1,\dots,T$ do
2:  if $t>0$ then
3:   Set $Q_{g}^{t}\leftarrow Q_{g}^{t-1}$
4:   Initialize new item embeddings: $Q_{g,i}^{t}$ for $i\in\mathcal{I}^{t+1}\setminus\mathcal{I}^{t}$
5:   Initialize new users: $\Phi_{u}$ for $u\in\mathcal{U}^{t+1}\setminus\mathcal{U}^{t}$
6:  end if
7:  for each round $r=1,2,\dots,R$ do
8:   $\mathcal{U}^{t,r}$ $\leftarrow$ Server randomly selects a subset of clients
9:   for each client $u\in\mathcal{U}^{t,r}$ in parallel do
10:     $Q_{u}^{t,r}\leftarrow\texttt{ClientUpdate}(u,Q_{g}^{t,r},t)$
11:   end for
12:   Compute $Q_{g^{\prime}}^{t,r}\leftarrow\frac{1}{|\mathcal{U}^{t,r}|}\sum_{u\in\mathcal{U}^{t,r}}Q_{u}^{t,r}$
13:   if $t=0$ then
14:     Set $Q_{g}^{t,r+1}\leftarrow Q_{g^{\prime}}^{t,r}$
15:   else
16:     Obtain $Q_{g}^{t,r+1}$ using Eq.13
17:   end if
18:  end for
19:end for
@@ 5.2. Main Performance
Misalignment of existing CLRec with FedRec. Table 2 shows that F3CRec outperforms RLP2 and SPP, which adapt centralized CLRec methods to FedRec. F3CRec surpasses RLP2 by 4% to 22%. While RLP2 leverages global knowledge for modeling preference shifts, FL constraints limit it to local knowledge, leading to performance drops. In contrast, F3CRec captures preference shifts individually and item knowledge shifts globally, yielding superior performance. SPP achieves the best baseline performance in about half of the cases, yet F3CRec still outperforms SPP by 2% to 16%. Unlike SPP’s uniform retention across users and items, F3CRec uses personalized retention at the client and item-specific weighting at the server, enhancing effectiveness. F3CRec also outperforms Reg and KD, which rely solely on user-specific data, with gains of 4%–50% and 1%–33%, respectively. Reg applies a naive retention approach without considering preferences, whereas KD assumes equal retention for all users. F3CRec explicitly accounts for individual preference shifts, leading to superior results. Overall, existing CLRec methods are misaligned with FedRec due to neglecting decentralization. F3CRec effectively addresses this by retaining knowledge on both client and server sides.
@@ 5.3.2. Analysis on Adaptive Replay Memory
ther the proposed preference shift measure (Sec 4.2.1) effectively distinguishes static users and dynamic users. We calculate the preference shift (Eq.2) for users with histories in both $\mathcal{D}^{0}$ and $\mathcal{D}^{1}$ after the first training round on $\mathcal{D}^{1}$. Users in the bottom 20% of preference shift are labeled static, and those in the top 20% dynamic. We then evaluate preference changes. The model trained on $\mathcal{D}^{0}$ is first evaluated on its test set of the same block. The model is then further trained on $\mathcal{D}^{1}$ and re-evaluated on the test set of $\mathcal{D}^{0}$ to assess knowledge preservation (Table 5) and performance degradation (Table 4), which is measured as $\frac{a_{0,0}-a_{1,0}}{a_{0,0}}$, where $a_{(t,s)}$ denotes performance when trained on $\mathcal{D}^{t}$ and evaluated on $\mathcal{D}^{s}$. The experiments are conducted using FedMF with the N@20 metric on the ML-100K and Lastfm-2K datasets.
Results in Table 4 and 5 demonstrate that the proposed method effectively distinguishes static and dynamic users. Static users consistently achieve higher N@20 and lower preference degradation rates while dynamic users show the opposite trend. This demonstrates the effectiveness of the ranking discrepancy-based preference shift measure and highlights its capability for adaptive, user-specific continual learning leveraging solely users’ own data without others.
@@ 5.6. Privacy-Preserving
Although private parameters and interaction data remain private in FL settings, information leakage may occur when transmitting public parameters to the server. Given a user $u$ with $n_{u}^{+}$ positive and $n_{u}^{-}$ negative items, the server can compare $Q_{g}^{t,r-1}$ and $Q_{u}^{t,r}$ to estimate the probability of an item being positive as $n_{u}^{+}/(n_{u}^{+}+n_{u}^{-})$, compromising privacy. To mitigate this, we add zero-mean Laplace noise to user updates before transmission: $Q_{u}^{t,r}=Q_{u}^{t,r}+\text{Laplace}(0,\lambda)$, where $\lambda$ controls the noise intensity. While not a formal local differential privacy (LDP) mechanism, this strategy is loosely motivated by LDP principles (Zhang et al., 2024b, 2023; Choi et al., 2018). We evaluate the performance of F3CRec with FedMF under varying noise intensities $\lambda\in\{0,0.1,0.2,0.3,0.4,0.5\}$ across four datasets, averaged over all data blocks (Table 8). Results show that even with stronger noise, performance degradation remains low, indicating our method reduces privacy leakage during aggregation without sacrificing utility.